Given the equation in X, (m + 1)X^2+ 4(m – 1)X + 4 – m = 0, determine m in each of

the following cases:

a) The equation has a double root, what is the root? (3marks)

b) the equation is of the first degree. What is the root? (3marks)

c) 0 is a root of the equation. What is the other root? (3marks)

d) the equation has two roots of different signs (3marks)

e) the equation has two positive roots. (3marks)


1
Expert's answer
2021-09-20T15:47:59-0400

"(m+1)x^2+4(m-1)x+4-m=0"


a) The equation has the double root if "D=4^2(m-1)^2-4(m+1)(4-m)=20m^2-44m=20m(m-2.2)=0" .

If "m=0" , then we have "x^2-4x+4=(x-2)^2=0" . The root is "x=2" .

If "m=2.2" , then we have "3.2x^2+4.8x-1.8=3.2(x-0.75)^2=0" . The root is "x=0.75"

b) If the equation is of the first degree, then "m+1=0" and "m=-1" .

We have the following equation: "-8x+5=0"

The root is "x=5\/8=0.625"


c) If 0 is the root of the equation, then "4-m=0" and "m=4" .

We have the following equation: "5x^2+12x=0"

"5x^2+12x=5x(x+2.4)=0"

The other root is "x=-2.4"


d) The equation has two roots of different signs, if "D>0" (existence) and "x_1x_2<0" (different signs).

"\begin{cases}
D=22m(m-2.2)>0
\\
x_1x_2=\frac{c}{a}=\frac{4-m}{m+1}<0
\end{cases}"


"\begin{cases}
m\in (-\infty , \ 0)\cup (2.2,\ +\infty)
\\
m\in (-\infty,\ -1)\cup (4,\ +\infty)
\end{cases}"


"m\in (-\infty,\ -1)\cup (4,\ +\infty)"


e) The equation has two (distinct) postings roots, if "D>0" (existence) and "x_1x_2>0" , "x_1+x_2>0" (positive).

"\begin{cases}
D=22m(m-2.2)>0
\\
x_1x_2=\frac{c}{a}=\frac{4-m}{m+1}>0
\\
x_1+x_2=-\frac{b}{a}=-\frac{4(m-1)}{m+1}>0
\end{cases}"


"\begin{cases}
m\in (-\infty , \ 0)\cup (2.2,\ +\infty)
\\
m\in (-1,\ 4)\\
m\in (-1,\ 1)
\end{cases}"


"m\in(-1,0)"


Answers:

a) "m=0,\ x=2" and "m=2.2, \ x=0.75"

b) "m=-1,\ x=0.625"

c) "m=4,\ x=-2.4"

d) "m\in (-\infty,\ -1)\cup (4,\ +\infty)"

e) "m\in(-1,0)"